OpenStudy (anonymous):

The width of a rectangle is 16 feet less than 3 times the lenght. If the area is 35 square feet, find the dimensions of a rectangle.

OpenStudy (anonymous):

width = 3 x length - 16 w =3l-16 A = lw = l(3l-16) A = 3l^2-16l 35 = 3l^2 - 16l 3l^2 - 16l - 35 = 0 (3l + 5)(l-7) = 0 3l + 5 = 0 l = -5/3 (not a valid measurement, so not the solution) or l-7 = 0 and l = 7 length is 7 and width is 3(7)-16=5